OpenStudy (anonymous):

how to integrate 1/(xlnx)?

OpenStudy (anonymous):

i'm trying to do it in parts, u=lnx and dv=x^(-1)dx

OpenStudy (anonymous):

but then i can't get v

OpenStudy (anonymous):

Let u = ln(x) du = 1/x dx => dx = xdu.

OpenStudy (anonymous):

plug that in and ull get a simple int 1/u du =]

OpenStudy (anonymous):

oh that's not even parts, that's just a simple substitution. wow, didn't see that there! thank you!